# Generalizing Goodstein’s theorem and Cichon’s independence proof

### v1

@@ 2.2 Quotients
      Otherwise, setting $\xi_{1}:=\Gamma+\rho$, check whether there is a (shortest) sequence $\xi_{1},\ldots,\xi_{m+1}$ (where $m\geq 1$) that determines a term $\Delta$ with $\chi^{\Omega_{j+1}}(\Delta)=1$ in the first step, such that
@@
        $\xi_{k}$ is of a form $\Delta[\vartheta_{j}(\xi_{k+1})]$ according to the lemma’s conditions (here according to condition 2. with i=j) for $k=1,\dots,m$, and
@@
      $\Gamma$ is of a form $\Delta[\zeta]$ according to the lemma’s conditions, so that $\alpha=\vartheta_{j}(\Delta[\zeta]+\rho)$ where $\eta:=\eta(j,\Delta,\rho)$ is defined, so that $\beta=\vartheta_{j}(\Delta+\eta)$, $\underline{\beta}=\rho$, and $\alpha=\beta[\zeta]$.
@@
Note that the outcome in case 4 (b) might be that $m=1$ and $\alpha=\beta[1]$, so that $\alpha$ fails to match the form detected for. Considering the example $\varepsilon_{0}$ as in the introduction, $\varepsilon_{0}[2]=\omega^{\omega}$ is seen as the first iterative repetition, despite $\varepsilon_{0}[1]=\omega=\vartheta_{0}(\vartheta_{0}(0))$.
We are now going to show that the <<-order type of the set ${{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}={\nicefrac{{{\mathring{{\operatorname{T}}}}}}{{k}}}\cap\Omega_{1}$ (where $k\in[2,\omega)$) is $\omega$ and that $\mathrm{G}_{k}$} enumerates this set, preserving its <<-order. To this end, we first define the analogue $\mathrm{imc}$} (for iterative maximal coefficient) of the measure $\operatorname{mc}$} that was used to define quotients $\nicefrac{{{\mathcal{E}}}}{{k}}$} of $\mathcal{E}$}. Recall Definition 2.4.
@@
The iterative maximal coefficient $\mathrm{imc}$} of terms in $\operatorname{T}$} is defined as follows.
@@
    ${\mathrm{imc}}(0):=0$.
@@
    If $\alpha\in{\operatorname{T}}$ is of a form $\vartheta_{i}(\xi)$ then
    ${\mathrm{imc}}(\alpha):=\max\{{\mathrm{imc}}(\xi),{\mathrm{imc}}(\underline{\alpha}),n\},$
    where either $n\geq 2$ is maximal such that $\alpha$ is of a form $\alpha=\beta[n]$ where $\beta\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$, or otherwise $n=1$.
Note that we then have ${\mathrm{imc}}(\alpha)=0$ if and only if $\alpha=0$. $\mathrm{imc}$} is weakly monotone with respect to the subterm relation, hence weakly increasing along localization sequences. But it also needs to be monotone with respect to support terms (which as mentioned before need not be subterms). Consider the ordinals $\varepsilon_{0}\cdot\omega$ and $\omega^{\omega+1}$, where $\mathrm{G}_{2}$} yields the result $8$, but the former is in ${\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[2]$}}}$} while the latter is not.
@@
For $\alpha\in{\operatorname{T}}\cap\mathrm{Lim}$ and $\zeta\in{\operatorname{T}}\cap\aleph_{\mathrm{d}(\alpha)}$ we have
${\mathrm{imc}}(\alpha[\zeta])\leq\max\{{\mathrm{imc}}(\alpha),{\mathrm{imc}}(\zeta)\}.$
@@
In particular, for $\alpha\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$ we have ${\mathrm{imc}}(\alpha[n])\leq\max\{{\mathrm{imc}}(\alpha),n\}$.
Proof. The lemma is proved by induction along Definition 2.3. The i.h. immediately yields the claim for additively decomposable limit ordinals. Now suppose that $\alpha=\vartheta_{i}(\Delta+\eta)>1$.
@@
    If $\eta\in\mathrm{Lim}$ and $\neg F_{i}(\Delta,\eta)$, that is, $\eta\in\mathrm{Lim}\cap\sup_{\sigma<\eta}\vartheta_{i}(\Delta+\sigma)$, we have $d=d(\eta)$ and $\alpha[\zeta]=\vartheta_{i}(\Delta+\eta[\zeta])$, and the i.h. for $\eta$ applies. The term $\underline{\alpha[\zeta]}$ is either 0, a subterm of $\alpha[\zeta]$, or $\eta[\zeta]$ is a successor ordinal. In this latter case we must either have $\eta=\xi+\omega$ for some $\xi\in\mathrm{Lim}\cup\{0\}$ along with $\zeta\in(0,\omega)$, or $\eta=\xi+{\Omega_{j}}$ for some $\xi$ such that ${\Omega_{j}}\mid\xi$ and $j\leq i$ along with $\zeta$ a successor ordinal (represented in $\operatorname{T}$}), say $\zeta=\nu+l+1$ for some $\nu\in\mathrm{Lim}\cup\{0\}$ and $l<\omega$. Note that then $\alpha[\zeta]=\vartheta_{i}(\Delta+\xi+\nu+\omega)[l+1]$ and $l+1\leq{\mathrm{imc}}(\zeta)$. We obtain $\underline{\alpha[\zeta]}=\vartheta_{i}(\Delta+\xi+\zeta-1)$, so that ${\mathrm{imc}}(\underline{\alpha[\zeta]})\leq\max\{{\mathrm{imc}}(\alpha),{\mathrm{imc}}(\zeta)\}$.
@@
    If otherwise $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$, the definition of $\alpha[\zeta]$ distinguishes between the following 3 subcases.
    1. (a)
      If $\Delta=0$, we have
      $\alpha[\zeta]=\left\{\begin{array}[]{cl}\underline{\alpha}\cdot\zeta&\mbox{ if }\eta>0\mbox{ (and hence $d=0$)}\\ \zeta&\mbox{ otherwise,}\end{array}\right$.
      so that in case of $\eta=0$ the claim is obvious, and in case of $\eta>0$ we have $\alpha[\zeta]=\underline{\alpha}\cdot\zeta$ where $\zeta<\omega$.
    2. (b)
      $\chi^{\Omega_{i+1}}(\Delta)=1$. So we have $d=0$, $\zeta<\omega$, and according to the recursive definition
      $\alpha[0]:=\vartheta_{i}(\Delta[\underline{\alpha}])\quad\mbox{ and }\quad\alpha[n+1]:=\vartheta_{i}(\Delta[\alpha[n]])\mbox{ for }n<\omega.$
      We need to verify that ${\mathrm{imc}}(\underline{\alpha[\zeta]})\leq\max\{{\mathrm{imc}}(\alpha),{\mathrm{imc}}(\zeta)\}$, which immediately follows by side induction on $\zeta$ since in general $\Delta[\xi]$ cannot be a successor for $\xi$ either 0 or an additive principal number.
    3. (c)
      Otherwise. Then $d=d(\Delta)$ and
      $\alpha[\zeta]=\vartheta_{i}(\Delta[\zeta]+\underline{\alpha}).$
      Here $\Delta[\zeta]$ is a proper multiple of $\Omega_{i+1}$ since $\chi^{\Omega_{i+1}}(\Delta)=0$. Thus, the i.h. smoothly applies to $\Delta$ and we obtain the claim. $\Box$
Lemma 2.13
For $\lambda\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$ and $\beta\in{\mathring{{\operatorname{T}}}}\cap[\lambda[k],\lambda)$ where $k\geq 2$ we have ${\mathrm{imc}}(\beta)\geq k$ and hence
${\nicefrac{{{\mathring{{\operatorname{T}}}}}}{{k}}}\cap[\lambda[k],\lambda)=\emptyset.$
Proof. We first observe that by Lemma 2.12, for any $\beta\in{\mathring{{\operatorname{T}}}}$
${\mathrm{imc}}(\beta)<k\Rightarrow{\mathrm{imc}}(\beta[i])<k\mbox{ for any }i<k.$
Suppose now that for some $\lambda\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$, $\beta\in[\lambda[k],\lambda)$ is least such that ${\mathrm{imc}}(\beta)<k$. This implies that $\beta\in\mathrm{Lim}$ such that $\lambda[k]<\beta<\lambda$, as ${\mathrm{imc}}(\lambda[k])\geq k$ and clearly ${\mathrm{imc}}(\gamma)\leq{\mathrm{imc}}(\gamma+1)$ for any $\gamma$. Bachmann property yields $\lambda[k]\leq\beta[1]$, and since ${\mathrm{imc}}(\beta[1])<k$, we cannot have $\lambda[k]<\beta[1]$ because of the minimality of $\beta$. But we cannot have $\lambda[k]=\beta[1]$ either, as this would imply that ${\mathrm{imc}}(\beta[1])\geq k$. $\Box$
The following theorem is in correspondence with Lemma 1 and Remarks 1 and 2 of [1].
@@
Proof. This now follows by <<-induction on $\alpha\in{{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}$ via possibly iterated application of Lemma 2.13. $\Box$

### v2

@@ 2.1 Preliminaries
      Otherwise, setting $\xi_{1}:=\Gamma+\rho$, check whether there is a (shortest) sequence $\xi_{1},\ldots,\xi_{m+1}$ (where $m\geq 1$) that determines a term $\Delta$ with $\Omega_{j+1}\mid\Delta<\Omega_{j+2}$ and $\chi^{\Omega_{j+1}}(\Delta)=1$ in the first step, such that
@@
        $\xi_{k}$ is of a form $\Delta[\vartheta_{j}(\xi_{k+1})]$ which is according to the lemma’s conditions (here according to condition 2 with i=j) for $k=1,\dots,m$, and
@@
      $\Gamma$ is of a form $\Delta[\zeta]$ according to the lemma’s conditions where $\Omega_{j+1}\mid\Delta<\Omega_{j+2}$ and $\chi^{\Omega_{j+1}}(\Delta)=0$, so that $\alpha=\vartheta_{j}(\Delta[\zeta]+\rho)$, and $\eta:=\eta(j,\Delta,\rho)$ is defined, so that $\beta=\vartheta_{j}(\Delta+\eta)$, $\underline{\beta}=\rho$, and $\alpha=\beta[\zeta]$.
@@
Considering the example $\varepsilon_{0}$ as in the introduction, $\varepsilon_{0}[1]=\omega^{\omega}$ is the first iterative repetition, while $\varepsilon_{0}[0]=\omega=\vartheta_{0}(0)$.
2.2 Quotients
For any $\beta\in{\operatorname{T}}$ it can be syntactically detected whether $\beta$ is of a form $\beta=\alpha[n]$ for some $\alpha\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$ and $n\in[1,\omega)$. Before we establish this explicitly, we use it in the following
@@ 2.2 Quotients
    $0,1\in{\nicefrac{{{\operatorname{T}}}}{{k}}}$.
  2. 2.
  3. 3.
    If $\xi=_{\mathrm{\scriptscriptstyle{ANF}}}\xi_{1}+\ldots+\xi_{m}\in{\operatorname{T}}\cap\Omega_{i+2}$ where $m\geq 0$ and $i<\omega$, then $\alpha:=\vartheta_{i}(\xi)\in{\nicefrac{{{\operatorname{T}}}}{{k}}}$, provided that
    1. (a)
      ${\xi^{\prime}}\in{\nicefrac{{{\operatorname{T}}}}{{k}}}$ where ${\xi^{\prime}}:=1+\xi$ if $i=0$ and ${\xi^{\prime}}:=\xi$ otherwise,
    2. (b)
      $\alpha[0]\in{\nicefrac{{{\operatorname{T}}}}{{k}}}$,
    3. (c)
      $\alpha_{1},\ldots,\alpha_{m-1}\in{\nicefrac{{{\operatorname{T}}}}{{k}}}$ where $\alpha_{l}:=\vartheta_{i}(\xi_{1}+\ldots+\xi_{l})$ for $l=1,\ldots,m-1$ in case of $m>1$, and
    4. (d)
      $\alpha$ is not of a form $\beta[n]$ where $\beta\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$ and $n\geq k-1$.
[…]
Remark 2.13
  1. 1.
    Note the condition $1+j\in{\nicefrac{{{\operatorname{T}}}}{{k}}}$ for $\vartheta_{0}(j)=\omega^{1+j}\in{\nicefrac{{{\operatorname{T}}}}{{k}}}$ where $j<\omega$, which is why we introduced $\xi^{\prime}$}.
  2. 2.
    Note that any of these quotients $\nicefrac{{{\operatorname{T}}}}{{k}}$} and $\nicefrac{{{\mathring{{\operatorname{T}}}}}}{{k}}$} should be closed under the operations $\cdot[l]$ for l<k1l<k-1. Due to Bachmann property this only needs to be explicitly stipulated for $l=0$, as we shall see.
  3. 3.
    It is easy to see that we have
[…]
We are now going to show that the <<-order type of the set ${{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}={\nicefrac{{{\mathring{{\operatorname{T}}}}}}{{k}}}\cap\Omega_{1}$ (where $k\in[2,\omega)$) is $\omega$ and that $\mathrm{G}_{k}$} enumerates this set, preserving its <<-order. To this end, we first define the analogue $\mathrm{imc}$} (for iterative maximal coefficient) of the measure $\operatorname{mc}$} that was used to define quotients $\nicefrac{{{\mathcal{E}}}}{{k}}$} of $\mathcal{E}$}. Recall the notions $\operatorname{lv}$, $\operatorname{Par}_{i}$, and $\operatorname{ht}_{i}$} from Section 2.1.
Definition 2.14
The iterative maximal coefficient ${\mathrm{imc}}(\alpha)$ of terms $\alpha\in{\operatorname{T}}$ is defined (in case of $\alpha\in{\mathbb{P}}^{>1}$) by primary recursion along the $\operatorname{Par}_{i}(\alpha)$-subterm relation where $i:=\operatorname{lv}(\alpha)$, secondary recursion on ${\operatorname{ht}_{i}}(\alpha)$, and tertiary <<-recursion on $\alpha$ as follows.
  1. 1.
    ${\mathrm{imc}}(0):=0$ and ${\mathrm{imc}}(1):=1$.
@@
    If $\alpha\in{\operatorname{T}}$ is of a form $\vartheta_{i}(\xi)$ where $\xi=_{\mathrm{\scriptscriptstyle{ANF}}}\xi_{1}+\ldots+\xi_{m}$ then, setting
    • ${\xi^{\prime}}:=\left\{\begin{array}[]{cl}1+\xi&\mbox{ if }i=0\\ \xi&\mbox{ otherwise,}\end{array}\right$. and
    • $\alpha_{l}:=\vartheta_{i}(\xi_{1}+\ldots+\xi_{l})$ for $l=1,\ldots,m-1$,
    we define
    ${\mathrm{imc}}(\alpha):=\max\{{\mathrm{imc}}({\xi^{\prime}}),{\mathrm{imc}}(\alpha[0]),{\mathrm{imc}}(\alpha_{1}),\ldots,{\mathrm{imc}}(\alpha_{m-1}),n+1\},$
    where either $n\geq 1$ is maximal such that $\alpha$ is of a form $\alpha=\beta[n]$ where $\beta\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$, or otherwise $n=0$.
Remark 2.15
Note that we have ${\mathrm{imc}}(\alpha)=0$ if and only if $\alpha=0$. $\mathrm{imc}$} is weakly monotone with respect to the subterm relation, hence weakly increasing along localization sequences. Since for $\alpha\in{\operatorname{T}}\cap{\mathbb{P}}^{>1}$ the definition of ${\mathrm{imc}}(\alpha)$ depends on ${\mathrm{imc}}(\alpha[0])$, we could not simply define by recursion along (modified) build-up stages of $\alpha$. A straightforward induction along the build-up of $\alpha\in{\operatorname{T}}\cap{\mathbb{P}}^{>1}$ shows that for $j\leq i:=\operatorname{lv}(\alpha)$ we have $\operatorname{Par}_{j}(\alpha[0])\subseteq\overline{\operatorname{Par}_{j}(\alpha)}$ where $\overline{\>\cdot\>}$ denotes the closure under the operation $\cdot[0]$.
@@
For $\lambda\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$ and $\beta\in{\operatorname{T}}\cap[\lambda[k-1],\lambda)$ where $k\geq 2$ we have ${\mathrm{imc}}(\beta)\geq k$ and hence
${\nicefrac{{{\operatorname{T}}}}{{k}}}\cap[\lambda[k-1],\lambda)=\emptyset.$
Proof. We first observe that by definition of $\mathrm{imc}$}, for any $\beta\in{\operatorname{T}}$
${\mathrm{imc}}(\beta)<k\Rightarrow{\mathrm{imc}}(\beta[0])<k.$
Suppose now that for some $\lambda\in{\mathring{{\operatorname{T}}}}\cap\mathrm{Lim}$, there exists a least $\beta\in{\operatorname{T}}\cap[\lambda[k-1],\lambda)$ such that ${\mathrm{imc}}(\beta)<k$. As ${\mathrm{imc}}(\lambda[k-1])\geq k$, this implies that $\lambda[k-1]<\beta<\lambda$, with $\beta\in\mathrm{Lim}$ since clearly ${\mathrm{imc}}(\gamma)\leq{\mathrm{imc}}(\gamma+1)$ for any $\gamma$. Bachmann property yields $\lambda[k-1]\leq\beta[0]$, which however contradicts the minimality of $\beta$ since ${\mathrm{imc}}(\beta[0])<k$. $\Box$
The following Theorem 2.21 is in correspondence with Lemma 1 and Remarks 1 and 2 of [1]. Since the generalized Goodstein principle introduced here induces much faster growing operations on the natural numbers than just base-$k$ exponentiation, Lemma 1 of [1] does not generalize directly. It is therefore instructive to sketch an alternative proof for the fact that the mapping $\mathrm{G}_{k}$} restricted to the quotient $\nicefrac{{{\mathcal{E}}}}{{k}}$} is the Mostowski collapse of $\nicefrac{{{\mathcal{E}}}}{{k}}$} onto $\omega$. This proof, which follows from the next lemma, then does generalize to the quotients ${\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}$}. For limit $\alpha$, by $i_{\alpha}$ we count how many times we can iterate the operation $\cdot[k-1]$ while staying within $\mathrm{Lim}$.
Lemma 2.18
For $k\in[2,\omega)$ and $\alpha\in{\mathcal{E}}\cap\mathrm{Lim}$ let
$i_{\alpha}:=\max\left\{i<\omega\mid\forall j\leq i\>\left(\alpha[k-1]^{j}\in\mathrm{Lim}\right)\right\}$ (7)
where
$\alpha[l]^{0}:=\alpha\;\mbox{ and }\;\alpha[l]^{j+1}:=(\alpha[l]^{j})[l]\mbox{ for }l<\omega.$
The ordinal $\alpha[k-1]^{i_{\alpha}}$ then clearly is a successor-multiple of $\omega$, so that $\alpha_{0}:=\alpha[k-1]^{i_{\alpha}}-\omega$ is well-defined and
${\mathrm{G}_{k}}(\alpha)={\mathrm{G}_{k}}(\alpha[k-1]^{i_{\alpha}+1})={\mathrm{G}_{k}}(\alpha_{0})+k.$
Moreover, if $\alpha\in{\nicefrac{{{\mathcal{E}}}}{{k}}}\cap\mathrm{Lim}$ then
$\alpha_{0}\in{\nicefrac{{{\mathcal{E}}}}{{k}}}.$
Proof. This now follows by <<-induction on $\alpha\in{\nicefrac{{{\mathcal{E}}}}{{k}}}$. If $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta$, the claim follows directly from the i.h. for $\eta$. If $\alpha=\omega^{\beta+1}$ for some $\beta$, we have $\alpha_{0}=0$ in case of $\beta=0$. Otherwise we have $\alpha[k-1]=\omega^{\beta}\cdot(k-1)+\omega^{\beta}$, apply the i.h. to $\omega^{\beta}$, and the claim follows. If finally $\alpha=\omega^{\lambda}$ for some $\lambda\in\mathrm{Lim}$, by the i.h. for $\lambda$ we obtain $\lambda_{0}:=\lambda[k-1]^{i_{\lambda}}-\omega\in{\nicefrac{{{\mathcal{E}}}}{{k}}}$. In case of $\lambda_{0}=0$, we obtain the claim with $i_{\alpha}=i_{\lambda}+k$. Otherwise we have
$\alpha[k-1]^{i_{\lambda}+k+1}=\omega^{\lambda_{0}+k-1}\cdot(k-1)+\ldots+\omega^{\lambda_{0}}\cdot(k-1)+\omega^{\lambda_{0}}$
and obtain the claim with $i_{\alpha}=i_{\lambda}+k+1+i_{\omega^{\lambda_{0}}}$ after an application of the i.h. to $\omega^{\lambda_{0}}$. $\Box$
Corollary 2.19
For $k\geq 2$ the mapping $\mathrm{G}_{k}$} restricted to the quotient $\nicefrac{{{\mathcal{E}}}}{{k}}$} is the Mostowski collapse of $\nicefrac{{{\mathcal{E}}}}{{k}}$} onto $\mathbb{N}$}.
Proof. As in Lemma 2.17 for quotients $\nicefrac{{{\operatorname{T}}}}{{k}}$}, we have a corresponding situation in $\nicefrac{{{\mathcal{E}}}}{{k}}$}: For any $\lambda\in{\mathcal{E}}\cap\mathrm{Lim}$ and $\beta\in[\lambda[k-1],\lambda)$ we have ${\operatorname{mc}}(\beta)\geq k$, thus
${\nicefrac{{{\mathcal{E}}}}{{k}}}\cap[\lambda[k-1],\lambda)=\emptyset.$
By Lemma 2.18, for any $\alpha\in{\nicefrac{{{\mathcal{E}}}}{{k}}}\cap\mathrm{Lim}$ the mapping $\mathrm{G}_{k}$} counts the elements of
$\alpha_{0},\alpha]\cap{\nicefrac{{{\mathcal{E}}}}{{k}}}=\{\alpha_{0}\}\cup\{\alpha[k-1]^{i_{\alpha}}[j]\mid j<k-1\}\cup\{\alpha\}$
starting from ${\mathrm{G}_{k}}(\alpha_{0})$ to ${\mathrm{G}_{k}}(\alpha)$ in increasing order. $\Box$
We now transfer this argumentation to terms in $\operatorname{T}$}. Recall that ${\mathrm{end}}(\gamma)$ of an ordinal $\gamma$ denotes the last additive component of $\gamma$. Clearly, as seen by straightforward <<-induction on $\alpha$, iterated application of the operation $\cdot[j]$, $j<\omega$, to any $\alpha\in{\operatorname{T}}\cap{\mathbb{P}}^{>1}$ eventually reaches a successor-multiple of $\Omega_{l}$} where $l:=\operatorname{lv}(\alpha)$.
Lemma 2.20
Let $k\in[2,\omega)$, $\alpha\in{\nicefrac{{{\mathring{{\operatorname{T}}}}}}{{k}}}\cap\mathrm{Lim}$, $l:=\operatorname{lv}({\mathrm{end}}(\alpha))$, and define $i_{\alpha}$, extending (7), by
$i_{\alpha}:=\left\{\begin{array}[]{cl}\max\left\{i<\omega\mid\forall j\leq i\>\left(\alpha[k-1]^{j}\in\mathrm{Lim}\right)\right\}&\mbox{ if }l=0\\[5.69054pt] \min\left\{i<\omega\mid\mathrm{d}(\alpha[k-1]^{i})>0\right\}&\mbox{ if }l>0,\end{array}\right$. (8)
and $\alpha_{0}$ by
$\alpha_{0}:=\left\{\begin{array}[]{cl}\alpha[k-1]^{i_{\alpha}}-\omega&\mbox{ if }l=0\\[5.69054pt] \alpha[k-1]^{i_{\alpha}}&\mbox{ if }l>0.\end{array}\right$.
Then,
@@
    if $l>0$, for any $\beta\in{\nicefrac{{{\mathring{{\operatorname{T}}}}}}{{k}}}\cap\aleph_{d}$ where $d:=\mathrm{d}(\alpha[k-1]^{i_{\alpha}})$, we have $\chi^{{\Omega_{d}}}(\alpha_{0})=1$ and
    $\alpha_{0}[\beta]\in{\nicefrac{{{\mathring{{\operatorname{T}}}}}}{{k}}},$
@@
    if $l=0$, i.e. $\alpha\in{{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}}$, we have
    $\alpha_{0}\in{{\mathring{{\operatorname{T}}}}{\mbox{\raisebox{0.77496pt}{$\scriptstyle[k]$}}}},$
    and it follows that
    ${\mathrm{G}_{k}}(\alpha)={\mathrm{G}_{k}}(\alpha[k-1]^{i_{\alpha}+1})={\mathrm{G}_{k}}(\alpha_{0})+k.$
@@
Proof. Using Lemmas 2.17 and 2.20 we argue as in the proof of Corollary 2.19. $\Box$